# Unraveling Zeroth-Order Optimization through the Lens of Low-Dimensional Structured Perturbations

### v1

@@ Abstract
Zeroth-order (ZO) optimization has emerged as a promising alternative to gradient-based backpropagation methods, particularly for black-box optimization and large language model (LLM) fine-tuning. However, ZO methods suffer from slow convergence due to high-variance stochastic gradient estimators. While structured perturbations, such as sparsity and low-rank constraints, have been explored to mitigate these issues, their effectiveness remains highly under-explored. In this work, we develop a unified theoretical framework that analyzes both the convergence and generalization properties of ZO optimization under structured perturbations. We show that high dimensionality is the primary bottleneck and introduce the notions of stable rank and effective overlap to explain how structured perturbations reduce gradient noise and accelerate convergence. Using the uniform stability under our framework, we then provide the first theoretical justification for why these perturbations enhance generalization. Additionally, through empirical analysis, we identify that block coordinate descent (BCD) to be an effective structured perturbation method. Extensive experiments show that, compared to existing alternatives, memory-efficient ZO (MeZO) with BCD (MeZO-BCD) can provide improved converge with a faster wall-clock time/iteration by up to ×$\times\textbf{2.09}$ while yielding similar or better accuracy.
Machine Learning, ICML
@@ 1 Introduction
Zeroth-order (ZO) optimization (Shamir, 2013), sometimes referred to as gradient-free or derivative-free optimization, has emerged as a compelling alternative to traditional gradient-based methods in deep learning. Unlike first- or higher-order approaches that rely on an explicit gradient or Hessian information, ZO methods update model parameters solely based on the function evaluations. This property makes them particularly useful in settings where gradients are inaccessible or expensive to compute, such as black-box optimization (Cai et al., 2021; Zhang et al., 2022b), adversarial attacks (Chen et al., 2017; Kurakin et al., 2017; Papernot et al., 2017), hyperparameter tuning (Li et al., 2021), neural architecture search (Wang et al., 2022), and efficient machine learning (Zhang et al., 2024a).
@@
Beyond these traditional applications, recently ZO optimization has gained attention for fine-tuning large language models (LLMs) (Malladi et al., 2023; Guo et al., 2024; Gautam et al., 2024; Liu et al., 2024b; Zhang et al., 2024d; Chen et al., 2024; Yu et al., 2024; Park et al., 2024), primarily due to its significantly lower memory requirements compared to gradient-based optimizations. A pioneering study, MeZO (Malladi et al., 2023), demonstrated that zeroth-order stochastic gradient descent (SGD) (Robbins & Monro, 1951) can fine-tune language models using memory and computational resources comparable to inference, making it a viable alternative for resource-constrained settings.
@@
However, despite its practicality, ZO optimization suffers from notoriously slow convergence and poor generalization, particularly in large-scale fine-tuning scenarios. Unlike first-order methods that benefit from exact gradient updates, ZO approaches rely on stochastic gradient estimates obtained via random perturbations, leading to inherently high variance and inefficient updates. While recent studies have attempted to address this issue by introducing structured perturbations, such as sparse perturbations (Liu et al., 2024b; Guo et al., 2024) and low-rank perturbations (Chen et al., 2024; Yu et al., 2024), the underlying principles that govern their effectiveness remain under-explored. Specifically, no unified framework exists to explain why these structures improve ZO efficiency or which structural properties are critical. Furthermore, the generalization properties of ZO optimization also remain elusive, as existing theoretical studies primarily focus on convergence while neglecting its impact on model generalization.
In this work, we provide a unified framework to analyze both the convergence and generalization properties of ZO optimization. Our theoretical contributions are as follows:
  • We present a unified theoretical framework for ZO optimization that formally decomposes the noise in gradient estimates into three fundamental factors. Our analysis highlights that high dimensionality emerges as the primary bottleneck, significantly amplifying variance and thereby slowing convergence.
  • We present a novel theoretical perspective grounded on the stable rank of perturbations and the intrinsic dimensionality of the loss landscape, capturing their effective overlap. Our analysis demonstrates that encouraging suitable low-dimensional constraints can substantially reduce noise and accelerate convergence. Moreover, our framework provides a principled explanation for why existing approaches, such as sparse and low-rank perturbations, enhance the efficiency of zeroth-order optimization.
  • Building on our unified framework, we establish a generalization error bound for zeroth-order optimization using the uniform stability. Our analysis shows that low-dimensional constraints also matter for generalization.
Leveraging these theoretical foundations, we conduct a comprehensive empirical study and uncover the potential of block coordinate descent (BCD) as an efficient alternative for ZO optimization. Experiments on LLM fine-tuning further validate the efficacy of MeZO with BCD (namely, MeZO-BCD). In specific, compared to existing alternatives, MeZO-BCD can provide improved converge with a faster wall-clock time per iteration by up to ×$\times\textbf{2.09}$ while yielding similar or better accuracy. These findings establish it as a promising method for large-scale ZO optimization.
@@ 2 Preliminaries
2.1 Notations
In this paper, we denote vectors by boldface lowercase letters (e.g. $\mathbf{x},\mathbf{y}$) and matrices by boldface uppercase letters (e.g. $\mathbf{A},\mathbf{B}$). We let $\mathcal{L}(\bm{\theta})$ denote the loss function with respect to $\bm{\theta}=(\theta_{1},\cdots,\theta_{d})\in\mathbb{R}^{d}$. Further, for any $\mu>0$, we define $\mathcal{L}_{\mu}(\bm{\theta})=\mathbb{E}_{\mathbf{u}}\big{[}\mathcal{L}(\bm{\theta}+\mu\mathbf{u})\big{]}$, where $\mathbf{u}\sim\mathcal{N}(\bm{0},\mathbf{I}_{d})$. For a matrix $\mathbf{A}$, we let $\sigma_{i}(\mathbf{A})$/$\lambda_{i}(\mathbf{A})$ denote $i$-th singular value/eigenvalue respectively, and $\sigma_{\text{max}}(\mathbf{A})$/$\lambda_{\text{max}}(\mathbf{A})$ represent the maximum singular value/eigenvalue of $\mathbf{A}$. Finally, we use $|\!|\!|\mathbf{A}|\!|\!|_{p}$ to denote the matrix $p$-norm.
2.2 Zeroth-Order Optimization
Simultaneous Perturbation Stochastic Approximation (SPSA). In this paper, we consider the following zeroth-order gradient estimator known to be SPSA as
@@ 2.2 Zeroth-Order Optimization
where $\mu$ is the smoothing parameter, $\mathbf{u}\sim\mathcal{N}(\bm{0},\mathbf{I}_{d})$, and $\mathcal{B}_{t}$ denotes minibatch selected at time $t$. Importantly, the SPSA estimator is unbiased estimator of $\nabla\mathcal{L}_{\mu}(\bm{\theta})$.
@@
3 Importance of Low-Dimensional Constraint on Convergence/Generalization
In this section, we provide the theoretical importance of low-dimensional constraints on convergence and generalization for zeroth-order optimization algorithms. Toward this, we first assume standard conditions.
@@ 3 Importance of Low-Dimensional Constraint on Convergence/Generalization
The condition (C-$1$) is standard in non-convex optimization analysis. In condition (C-$2$), the unbiasedness of stochastic gradients and bounded variance are fundamental in stochastic optimization literature (Ghadimi & Lan, 2013). The bounded gradient condition in (C-$2$) is frequently used in convergence analysis in the context of adaptive gradient methods (Kingma & Ba, 2015; Reddi et al., 2018; Chen et al., 2019a; Yun et al., 2022; Ahn & Cutkosky, 2024).
Under these assumptions, we start by introducing an important quantity, the zeroth-order gradient noise.
Definition 3.1 (Zeroth-order Gradient Noise).
Let $\widehat{\mathbf{g}}$ be any zeroth-order gradient estimator of first-order gradient $\nabla\mathcal{L}(\bm{\theta})$. Then, the ZO gradient noise of $\widehat{\mathbf{g}}$ is defined as
$\zeta_{\text{zo}}=\mathbb{E}\big{[}\|\widehat{\mathbf{g}}-\nabla\mathcal{L}(\bm{\theta})\|^{2}\big{]}.$
In fact, if $\widehat{\mathbf{g}}$ is the first-order stochastic gradient, the quantity $\zeta_{\text{zo}}$ just boils down to the traditional variance. However, in general, ZO gradient estimators are not unbiased, thus in order to avoid confusion, we call the quantity $\zeta_{\text{zo}}$ as zeroth-order gradient noise.
3.1 Cause of Zeroth-order Gradient Noise
Traditionally, ZO optimization algorithms suffer from notoriously slow convergence, specifically for training models on high-dimensional parameter spaces. The slow convergence of ZO optimization comes from the high noise of ZO gradient. In order to see for ZO gradient how far from the first-order gradient, we revisit the zeroth-order gradient evaluated on the single datapoint $\mathbf{z}$, which is computed using (1). The following lemma shows the zeroth-order gradient noise for the estimator in (1).
Lemma 3.2.
The zeroth-order gradient noise of $\mathbf{g}_{t}$ by
$\zeta_{\textnormal{zo}}\leq\underbrace{\frac{\mu^{2}L^{2}(d+6)^{3}}{2}}_{\textnormal{Smoothing Error}}+4\underbrace{(d+2)}_{\begin{subarray}{c}\textnormal{Dimension}\\ \textnormal{Error}\end{subarray}}\Big{(}G^{2}+\underbrace{\frac{\sigma^{2}}{B}}_{\begin{subarray}{c}\textnormal{Minibatch}\\ \textnormal{error}\end{subarray}}\Big{)}$
Note that the smoothing error could be arbitrarily small by controlling the parameter $\mu$ and the minibatch noise $\sigma^{2}/B$ is also inherent from the first-order gradient estimation. Thus, the dominant factor of zeroth-order noise comes from the high dimensionality of the parameter. In the next section, we illustrate how this zeroth-order noise could be reduced.
3.2 Reducing the Dimension Error: Perturbation Perspective
As described in Section 3.1, the most problematic part in zeroth-order gradient noise is the dimension error. Thus, to reduce the dependency of the parameter dimension, we investigate how the zeroth-order noise could be handled. Toward this, we consider the projected zeroth-order gradient in the sense that it is induced by some matrix $\mathbf{M}_{t}\in\mathbb{R}^{d\times d}$ at time $t$ as
$\widehat{\mathbf{g}}_{t}$ $=\frac{\mathcal{L}(\bm{\theta}_{t}+\mu\mathbf{M}_{t}\mathbf{u}_{t};\mathcal{B}_{t})-\mathcal{L}(\bm{\theta}-\mu\mathbf{M}_{t}\mathbf{u}_{t};\mathcal{B}_{t})}{2\mu}\mathbf{M}_{t}\mathbf{u}_{t}.$ (2)
In this definition, we assume for simplicity that $\mathbf{M}_{t}$ is a non-zero, symmetric positive semi-definite (PSD) matrix. We further adopt the mild condition that its largest singular value is bounded by some constant $C$, independent of the parameter dimension $d$. Without loss of generality, we set $C=1$ throughout this paper, since the singular values of $\mathbf{M}_{t}$ can be controlled by the smoothing parameter $\mu$.
With this projected gradient estimator in (2), we consider the following ZO-SGD
$\bm{\theta}_{t+1}=\bm{\theta}_{t}-\alpha_{t}\widehat{\mathbf{g}}_{t}.$ (3)
We introduce the stable rank (Ipsen & Saibaba, 2024) which is crucial for our arguments.
Definition 3.3 (Stable Rank).
For any matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, the stable rank $\mathrm{srank}(\mathbf{A})$ is defined by
$\mathrm{srank}(\mathbf{A})=\frac{\sum_{i}\sigma_{i}(\mathbf{A})^{2}}{\sigma_{\text{max}}(\mathbf{A})^{2}}.$
The stable rank of a matrix $\mathbf{A}$ is a robust measure of its effective dimensionality. Intuitively, whereas the usual (discrete) rank simply counts the number of non-zero singular values, the stable rank accounts for how large or small each singular value is, making it less sensitive to small perturbations. As a result, the stable rank provides a smoother estimate of the “spread” or “concentration” of the matrix’s singular values than the classical rank.
Now, we introduce our key assumption.
Assumption 3.4 (Small Stable Rank).
The stable rank of $\mathbf{M}_{t}$ satisfies
$\mathrm{srank}(\mathbf{M}_{t})\leq s\ll d$
for some constant $s=\Theta(1)$ with respect to $d$.
Imposing Assumption 3.4 provides a robust and flexible alternative to traditional low-rank constraints. Unlike the discrete rank, which can jump abruptly in response to small changes in singular values, the stable rank varies smoothly, thus more accurately reflecting the “effective” dimensionality. Furthermore, even if a matrix is formally full-rank, enforcing Assumption 3.4 can effectively capture near-sparse or near-low-rank structure by considering the distribution of singular values rather than simply counting their non-zero entries. Consequently, maintaining a small stable rank encourages approximate low-dimensional behavior across a broader range of real-world scenarios than the classical low-rank constraint. Note also that it always follows $1\leq\mathrm{srank}(\mathbf{M}_{t})\leq\mathrm{rank}(\mathbf{M}_{t})\leq d$ for non-zero $\mathbf{M}_{t}$.
Under this assumption, we can improve the zeroth-order gradient noise with the gradient estimator in (2).
Lemma 3.5.
Under (C-$1$), (C-$2$), and Assumption 3.4, the zeroth-order gradient noise of $\widehat{\mathbf{g}}_{t}$ in (2) is bounded by
$\zeta_{\textnormal{zo}}\leq\underbrace{\frac{\mu^{2}L^{2}(d+6)^{3}}{2}}_{\textnormal{Smoothing Error}}+4\underbrace{(s+2)}_{\begin{subarray}{c}\textnormal{Dimension}\\ \textnormal{Error}\end{subarray}}\Big{(}G^{2}+\underbrace{\frac{\sigma^{2}}{B}}_{\begin{subarray}{c}\textnormal{Minibatch}\\ \textnormal{Error}\end{subarray}}\Big{)}$
where $B$ denotes the batch size.
It can be clearly seen that the dimension error is dramatically reduced to the order of $s$. We further provide the convergence of ZO-SGD under small stable rank.
Theorem 3.6 (Convergence of ZO-SGD).
Suppose that (C-$1$), (C-$2$), and Assumption 3.4 hold. Under the following parameter settings
$\mu=\mathcal{O}\Big{(}\frac{1}{L\sqrt{sdT}}\Big{)},\quad\alpha\leq\frac{1}{8L(s+2)},$
ZO-SGD in (3) satisfies
$\mathbb{E}\big{[}\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}\big{]}\leq\mathcal{O}\Big{(}\underbrace{\frac{d^{2}}{sT}}_{\begin{subarray}{c}\textnormal{Smoothing}\\ \textnormal{Error}\end{subarray}}+\underbrace{\frac{s^{2}}{dT}}_{\begin{subarray}{c}\textnormal{Dimension}\\ \textnormal{Error}\end{subarray}}+\frac{\Delta}{\alpha T}+\underbrace{\frac{\sigma^{2}}{B}}_{\begin{subarray}{c}\textnormal{Minibatch}\\ \textnormal{Error}\end{subarray}}\Big{)},$
We make several remarks on Theorem 3.6.
On Convergence Results.  Under the small stable rank assumption, the dimension error in Theorem 3.6 can be bounded independently of $d$. Since $s$ is a parameter we can freely choose, adjusting $s$ affects the rate of convergence. In particular, as $s$ approaches $d$, the convergence rate becomes faster (smoothing error is dominant), eventually reducing to the vanilla ZO-SGD case (where $\mathbf{M}_{t}$ is the identity matrix), thus recovering the well-known complexity such as $\mathcal{O}(d/T)$. Note that the smoothing error is governed by the structure of the loss landscape (i.e. problem structure), thus cannot be reduced by any particular choice of $\mathbf{M}_{t}$. We will delve more deeply into the discussion for this issue in Section 3.3.
On Parameter Condition.  Regarding the smoothing parameter $\mu$, under Assumption 3.4, the order of $\mu$ approximately becomes $\mathcal{O}(\frac{1}{L\sqrt{dT}})$ which is very standard in zeroth-order optimization literature (Liu et al., 2018; Chen et al., 2019b). Note also that we could dramatically improve the condition on the learning rate. While the conventional choice in ZO optimization, such as $\alpha=\mathcal{O}(1/d)$, results in an impractically small learning rate, so small that training hardly progresses. We drastically relax this requirement to $\alpha=\mathcal{O}(1/s)$, thereby making the learning process far more practical.
We present two examples of how matrix $\mathbf{M}_{t}$ enforce a low-dimensional constraint: (i) sparsity and (ii) low-rankness.
Sparsity.  One of the simplest ways to encourage a low-dimensional constraint is to sparsify some portion of diagonal entries of the identity matrix to zero, thereby forming a matrix $\mathbf{M}_{t}$ of the following structure:
$\mathbf{M}_{\text{sparsity}}=\mathrm{diag}(\delta_{1},\delta_{2},\cdots,\delta_{d}),$
where $\delta_{i}\in\{0,1\}$ and $\sum_{i}\delta_{i}=s\ll d$. By applying this transformation to the perturbation, only a subset of parameters is updated at each iteration, rather than all parameters. This idea aligns with the spirit of recent studies (Liu et al., 2024b; Guo et al., 2024) that leverage sparsity in zeroth-order optimization for fine-tuning LLMs.
Low-Rankness. Another way to encourage a low-dimensional constraint is to exploit low-rankness. For instance, suppose that matrix $\mathbf{M}_{t}$ has the following form:
$\mathbf{M}_{\text{low-rank}}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$
where $\mathbf{U}\in\mathbb{R}^{d\times s}$ and $\mathbf{V}\in\mathbb{R}^{d\times s}$ are (random) orthonormal matrices and $\bm{\Sigma}\in\mathbb{R}^{s\times s}$ is diagonal. It is clear that the stable rank of $\mathbf{M}_{\text{low-rank}}$ is at most $s$, and according to our argument, using $\mathbf{M}_{\text{low-rank}}$ effectively reduces the dimension error. This approach has recently drawn increasing attention in zeroth-order fine-tuning of LLMs, and some studies (Rando et al., 2023; Yu et al., 2024; Chen et al., 2024) can be partially explained by our framework.
Note.  Although we showcase examples that exploit exact sparsity and exact low-rankness, our argument only requires that the stable rank of $\mathbf{M}_{t}$ be small. In other words, our theory naturally covers near-sparse or near-low-rank $\mathbf{M}_{t}$, thereby accommodating a much broader range of transformations. However, the smoothing error still depends on $d$ as shown in Theorem 3.6. In the next section, we discuss how to mitigate this issue from the loss perspective.
3.3 Towards Improved Convergence: Loss Perspective
Although we successfully reduce the dimension error in Theorem 3.6, the smoothing error is still problematic. In this section, we investigate how to alleviate the smoothing error in the perspective of loss landscapes. Toward this, we introduce the intrinsic dimension (Ipsen & Saibaba, 2024) of a matrix.
Definition 3.7 (Intrinsic Dimension).
For matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, the intrinsic dimension denoted by $\mathrm{intdim}(\mathbf{A})$ is defined by
$\mathrm{intdim}(\mathbf{A})=\frac{\sum_{i}\lambda_{i}(\mathbf{A})}{\lambda_{\text{max}}(\mathbf{A})}.$
Using the intrinsic dimension, Malladi et al. (2023) propose an assumption that the local Hessian of loss landscape has a low intrinsic dimension (in fact, the authors use the term “local effective rank”, but in this work, to avoid the confusion with the stable rank, we use the terminology an intrinsic dimension). However, Malladi et al. (2023) only consider the limit case of $\mu\rightarrow 0$, thus it is not realistic for practical scenario since the zeroth-order optimization algorithms actually employ strictly positive $\mu>0$. Also, Malladi et al. (2023) consider the original SPSA in (1), which suffers from the extremely high noise as we discuss in Lemma 3.2. Therefore, in this work, we propose the revised condition that is also compatible with our Assumption 3.4.
Assumption 3.8 (Locally Low Intrinsic Dimension).

### v2

# Elucidating Subspace Perturbation in Zeroth-Order Optimization: Theory and Practice at Scale

@@ Abstract
Zeroth-order (ZO) optimization has emerged as a promising alternative to gradient-based backpropagation methods, particularly for black-box optimization and large language model (LLM) fine-tuning. However, ZO methods often suffer from slow convergence due to high-variance stochastic gradient estimators. While subspace perturbations, such as sparsity and low-rank constraints, have been explored to mitigate this issue, their effectiveness remains poorly understood. In this work, we develop a unified theoretical framework that analyzes both the convergence and generalization properties of ZO optimization under subspace perturbations. We show that high dimensionality is the primary bottleneck and introduce the notion of subspace alignment to explain how the subspace perturbations reduce gradient noise and accelerate convergence. Our analysis further shows that a broad class of subspace perturbations exhibits a similar convergence rate, motivating us to prioritize practical considerations in real-world algorithm design. Building on these insights, we propose an efficient ZO method using block coordinate descent (MeZO-BCD), which perturbs and updates only a subset of parameters at each step. Extensive experiments show that MeZO-BCD significantly accelerates optimization, achieving up to ×$\mathbf{\times 2.77}$ speedup in wall-clock time over MeZO on OPT-13B, while maintaining comparable iteration complexity and fine-tuning performance.
@@ 1 Introduction
Zeroth-order (ZO) optimization (Shamir, 2013), sometimes referred to as gradient-free or derivative-free optimization, has emerged as a compelling alternative to traditional gradient-based methods in deep learning. Unlike first- or higher-order approaches that rely on explicit gradient or Hessian information, ZO methods update model parameters solely based on the function evaluations. This property makes them particularly useful in settings where gradients are inaccessible or expensive to compute, such as black-box optimization (Cai et al., 2021; Zhang et al., 2022b), adversarial attacks (Chen et al., 2017; Kurakin et al., 2017; Papernot et al., 2017), hyperparameter tuning (Li et al., 2021), neural architecture search (Wang et al., 2022), and efficient machine learning (Zhang et al., 2024a).
@@
Beyond these traditional applications, recently ZO optimization has gained attention for fine-tuning large language models (LLMs) (Malladi et al., 2023; Guo et al., 2024; Gautam et al., 2024; Liu et al., 2024b; Zhang et al., 2024e; Chen et al., 2024; Yu et al., 2024; Park et al., 2024), primarily due to its significantly lower memory requirements compared to gradient-based optimizations. A pioneering study, MeZO (Malladi et al., 2023), demonstrated that zeroth-order stochastic gradient descent (SGD) (Robbins and Monro, 1951) can fine-tune language models using memory and computational resources comparable to inference, making it a viable alternative for resource-constrained settings.
@@
However, despite its practicality, ZO optimization suffers from notoriously slow convergence and poor generalization, particularly in large-scale fine-tuning scenarios. Unlike first-order methods that benefit from exact gradient updates, ZO approaches rely on stochastic gradient estimates obtained via random perturbations, leading to inherently high variance and inefficient updates. While recent studies have attempted to address this issue by introducing subspace perturbations, such as sparse perturbations (Liu et al., 2024b; Guo et al., 2024) and low-rank perturbations (Chen et al., 2024; Yu et al., 2024), the underlying principles that govern their effectiveness remain under-explored. Specifically, no unified framework exists to explain why these structures improve ZO efficiency or which structural properties are critical. Furthermore, the generalization properties of ZO optimization also remain elusive, as existing theoretical studies primarily focus on convergence while neglecting its impact on model generalization. In light of these observations, it is natural to pose the following question:
“What makes subspace perturbation effective, and how can this guide better algorithms?”
In this paper, we answer this question by providing a unified framework to analyze both the convergence and generalization properties of ZO optimization. This analysis shows that diverse perturbation strategies exhibit similar convergence behavior despite their structural differences, motivating algorithm design driven by practical efficiency rather than subspace structure. Building on this insight, we develop a practical ZO method for large-scale optimization. Our main contributions can be summarized as follows:
  • Unified theoretical framework for subspace perturbation. We introduce a unified analysis of convergence and generalization in ZO optimization under subspace perturbations, based on the notions of subspace alignment and dimensionality. Importantly, our analysis remains valid under finite perturbation magnitude, avoiding the restrictive infinitesimal smoothing assumption commonly adopted in prior works, hence enabling more realistic modeling.
  • Theoretical criteria for effective subspace perturbation. We show that sparse, low-rank, and other common structures share the same expected subspace alignment, implying equivalent average-case convergence. This enables the design of ZO algorithms that prioritize practical considerations, such as computational and memory efficiency, when choosing subspace perturbation.
  • MeZO-BCD: a practical and scalable ZO algorithm. Based on our theory, we propose MeZO-BCD, a block coordinate descent method for ZO that perturbs and updates only one block per step. This substantially reduces memory transfer overhead and achieves up to ×$\mathbf{\times 2.77}$ wall-clock speedup over MeZO on OPT-13B, while maintaining convergence and fine-tuning performance. The approach also enables promising future extensions such as adaptive block selection and integration with stateful optimizers.
@@ 2 Preliminaries
Notations. In this paper, we denote vectors by boldface lowercase letters (e.g. $\mathbf{x},\mathbf{y}$) and matrices by boldface uppercase letters (e.g. $\mathbf{A},\mathbf{B}$). We let $\mathcal{L}(\bm{\theta})$ denote the loss function with respect to $\bm{\theta}=(\theta_{1},\cdots,\theta_{d})\in\mathbb{R}^{d}$. Further, for any $\mu>0$, we define $\mathcal{L}_{\mu}(\bm{\theta})=\mathbb{E}_{\mathbf{u}}\big{[}\mathcal{L}(\bm{\theta}+\mu\mathbf{u})\big{]}$, where $\mathbf{u}\sim\mathcal{N}(\bm{0},\mathbf{I}_{d})$. For a matrix $\mathbf{A}$, we let $\sigma_{i}(\mathbf{A})$/$\lambda_{i}(\mathbf{A})$ denote $i$-th singular value/eigenvalue respectively, and $\sigma_{\text{max}}(\mathbf{A})$/$\lambda_{\text{max}}(\mathbf{A})$ represent the maximum singular value/eigenvalue of $\mathbf{A}$. Also, we use $|\!|\!|\mathbf{A}|\!|\!|_{p}$ to denote the matrix $p$-norm. Finally, we define two quantities, stable rank and intrinsic dimension of matrix: $\mathrm{srank}(\mathbf{A})\coloneqq\frac{\sum_{i}\sigma_{i}(\mathbf{A})^{2}}{\sigma_{\text{max}}(\mathbf{A})^{2}}$ and $\mathrm{intdim}(\mathbf{A})\coloneqq\frac{\mathrm{Tr}(\mathbf{A})}{|\!|\!|\mathbf{A}|\!|\!|_{\mathrm{op}}}$ (Ipsen and Saibaba, 2024).
Simultaneous Perturbation Stochastic Approximation (SPSA). We consider the following zeroth-order gradient estimator, known to be SPSA, as
@@
where $\mu$ is the smoothing parameter, $\mathbf{u}_{t}\sim\mathcal{N}(\bm{0},\mathbf{I}_{d})$, and $\mathcal{B}_{t}$ denotes minibatch selected at time $t$. Importantly, the SPSA estimator is an unbiased estimator of $\nabla\mathcal{L}_{\mu}(\bm{\theta})$.
@@
3 Theoretical Foundations on Subspace Perturbation
3.1 Subspace ZO-SGD and Standard Assumptions
We begin by formalizing the setup of zeroth-order optimization with structured perturbations. In particular, we describe the subspace ZO-SGD algorithm and introduce standard assumptions that underlie our convergence and generalization analysis.
@@ 3.1 Subspace ZO-SGD and Standard Assumptions
The condition (C-$1$) is standard in non-convex optimization analysis. In condition (C-$2$), the unbiasedness of stochastic gradients and bounded variance are fundamental in stochastic optimization literature (Ghadimi and Lan, 2013). The bounded gradient condition in (C-$2$) is frequently used in convergence analysis in the context of adaptive gradient methods (Kingma and Ba, 2015; Reddi et al., 2018; Chen et al., 2019; Yun et al., 2022; Ahn and Cutkosky, 2024).
Under these assumptions, we consider a ZO-SGD with subspace perturbation:
$\widehat{\mathbf{g}}_{t}$ $=\frac{\mathcal{L}(\bm{\theta}_{t}+\mu\mathbf{M}_{t}\mathbf{u}_{t};\mathcal{B}_{t})-\mathcal{L}(\bm{\theta}_{t}-\mu\mathbf{M}_{t}\mathbf{u}_{t};\mathcal{B}_{t})}{2\mu}\mathbf{M}_{t}\mathbf{u}_{t},$ (2)
where $\mu$ is called a smoothing parameter. For simplicity, we assume that $\mathbf{M}_{t}$ is positive semi-definite and its maximum singular value is $\sigma_{\text{max}}(\mathbf{M}_{t})=1$ since it can be controlled by $\mu$. Also, we refer to the zeroth-order SGD with subspace perturbation (2) as subspace ZO-SGD, i.e., the model parameter is updated via $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\alpha_{t}\widehat{\mathbf{g}}_{t}$.
3.2 Convergence Analysis of Subspace ZO-SGD
From this section, we provide the theoretical understanding of subspace ZO-SGD. Toward this, we start from convergence analysis. In recent studies, there have been many attempts to leverage the zeroth-order optimization in fine-tuning large language models due to its memory-efficiency (Malladi et al., 2023). To address both from-scratch training and fine-tuning scenarios, we propose the following assumption.
Assumption 3.1 (Local Intrinsic Dimension).